OpenStudy (anonymous):

i have done this problem multiple times and still cant get it. Help please (1 pt) The altitude of a triangle is increasing at a rate of 2.5 centimeters/minute while the area of the triangle is increasing at a rate of 1 square centimeters/minute. At what rate is the base of the triangle changing when the altitude is 9.5 centimeters and the area is 97?

OpenStudy (ranga):

A = 1/2 * b * h dA/dt = 1/2 * (b*dh/dt + h * db/dt) Plug in the numbers and find db/dt.

OpenStudy (anonymous):

I have done that, and im not getting the anwer. I get something like 5.424, every time which is wrong

OpenStudy (ranga):

A = 1/2 * b * h A = 97, h = 9.5 97 = 0.5 * b * 9.5 b = 20.42 dA/dt = 1/2 * (b*dh/dt + h * db/dt) dA/dt = 1 dh/dt = 2.5 b = 20.42 h = 9.5 1 = 0.5 * (20.42 * 2.5 + 9.5 * db/dt) 2 = 20.42 * 2.5 + 9.5 * db/dt 2 = 51.05 + 9.5 * db/dt db/dt = (2-51.05)/9.5 = -5.16 cm / minute.

OpenStudy (anonymous):

ok, I see what I did there. It was the area, I never took over the 1/2. Thank u!

OpenStudy (ranga):

You are welcome.